Fermi National Laboratory
 

Accelerator Update for April 25, 2007 - April 27, 2007

 

Wednesday April 25

The day shift began with Operations monitoring the Tevatron (TeV) store 5374 (142.45E30) and the Recycler stash (107.1E10), with the Antiproton Source (Pbar) stacking (8.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:16 AM, a Linac RF station (LRF3) tripped off. It reset.

A RF expert reported at 8:21 AM that MIRF9 needed its modulator changed and that work would take about 45 minutes.

A Switchyard (SY)/ Meson expert began tuning the MTest beam line at 8:34 AM.

At 11:26 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 11:36 AM, BRF3 tripped off. It reset.

At 11:37 AM, the I- Source vacuum gauge tripped off. It reset.

At 1 PM, Operations turned off the Linac beam for the following jobs: MIRF9 repair, tie in the MI-20L power supply, work on the Linac low level RF, and investigate the KRF4 problems.

Linac experts reported at 1:22 PM that a complete change out of the KRF4 Mod VME crate cards fixed the readback problems.

The Linac experts reported that they completed their work. Operations resumed sending beam to MiniBooNE at 1:59 PM.

RF experts reported that the MIRF9 work was completed. Operations resumed sending beam to Pbar, NuMI, and SY at 2:11 PM.

At 3:03 PM, MIRF14 tripped off. It reset.

At 3:07 PM, Operations reported that the beam had fallen out of the MI accelerator. An expert thought it was an RF problem. Operations rebooted the low level RF system and MI resumed operating at full beam intensity at 3:46 PM. (Also, the MI-20L power supply tie in wasn’t successful. MI continued to run on 11 power supplies.)

Operations reported at 3:59 PM that the Recycler had stashed antiprotons twice during the day shift.

SY experts finished using the beam at 5:35 PM. Operations disabled the beam.

At 5:38 PM, MIRF14 tripped off. It reset.

At 6:53 PM, MIRF14 tripped off again and would not turn on. Operations lowered MI’s beam intensity. 30 minutes later operators were able to reset the station and then brought the intensity back up to normal.

At 8:23 PM. LRF3 tripped off. It reset.

At 11:04 PM, KRF4 tripped off. It reset.

Operations reported at 11:55 PM that the Recycler had stashed four times during the evening shift.

Thursday April 26

The midnight shift began with Operations monitoring store 5374 (47.39E30) and the stash (293.6E10), with Pbar stacking (10.05E10), and with NuMI and MiniBooNE taking beam.

At 2:24 AM, the Pbar Debuncher Lithium Lens (D:LNV) tripped off due to a power supply high temperature. Operators left the power supply off so it could cool down, which took about 30 minutes. There was no stacking during the cool down.

At 4:10 AM, KRF4 tripped off. It reset.

At 7:19 AM, KRF1 tripped off for the first of two times in four minutes. It reset.

Operations terminated store 5374 at 8 AM.

Operations reported at 8:03 AM that the Recycler had stashed twice during the midnight shift.

Operations had CDF and D0 ready for access by 8:19 AM. Both experiments were out of access by 8:59 AM.

Operations established beam to MTest for experts to tune at 9 AM.

Operations put the TeV into a dry squeeze at 9:54 AM.

Operations put the TeV into shot setup at 10:23 AM.

Operations established store 5376 at 12:31 PM with an initial luminosity of 276.43E30.

At 2:08 PM, LRF3 tripped off. It reset.

At 2:37 PM, KRF3 tripped off. The RF cavity LCW valve stuck open and the cavity cooled, which changed its phase. Operators, with the help of an expert, managed to close the valve and then open it back up to an appropriate setting. At 3:10 PM, the temperature returned to normal, the cavity returned to its normal phase, and Linac resumed sending beam.

At 3:16 PM, LRF3 tripped off. It reset.

Operations reported at 4:08 PM that the Recycler had stashed twice during the day shift.

SY and MTest beam tuning ended 5:26 PM.

At 5:41 PM, KRF4 tripped off. It reset.

At 5:45 PM, BRF14 tripped off. It reset.

At 5:47 PM, LRF3 tripped off. It reset.

At 7:45 PM, TeV store 5376 aborted. There was no quench. The TeV sector A4 QPM had pulled the abort. Operations contacted experts.

Operations put the TeV through a dry squeeze at 9:22 PM.

Operations prepared CDF for an access at 9:53 PM. The experimenters came out of access at 11:55 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Friday April 27

The midnight shift began with Operations monitoring the stash (201.1E10), with Pbar stacking (31.7E10), and with NuMI and MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 12:21 AM.

Operations put the TeV into shot setup at 12:47 AM.

Operations established store 5379 at 3:03 AM with an initial luminosity of 156.39E30.

Operations reported at 8:05 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/27/2007    email Fermilab